Majer Corporation makes a product with the following standard costs: 






Standard Quantity or HoursStandard Price or  RateStandard Cost Per Unit
Direct materials6.2 ounces$5.00 per ounce$31.00
Direct labor0.6 hours$11.00 per hour$6.60
Variable overhead0.6 hours$5.00 per hour$3.00


The company reported the following results concerning this product in February.

Originally budgeted output5,500 units
Actual output5,600 units
Raw materials used in production33,000 ounces
Actual direct labor-hours2,020 hours
Purchases of raw materials35,400 ounces
Actual price of raw materials$7.10 per ounce
Actual direct labor rate$2.40 per hour
Actual variable overhead rate$5.20 per hour

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The variable overhead efficiency variance for February is:
Answer #1

Variable overhead efficiency variance

= (SLH - ALH) * SVOHR

= (5600*0.6 - 2020) * 5

= (3360 - 2020) * 5

= 6,700 Favourable

Comment if you face any issues